OpenStudy (anonymous):

one zero of f(x)=x^3-6x^2+11x-6 is 3 what are other zeros of the function A 1 and 3 B -1 and -2 C 1 and 2 D -1 and 3

OpenStudy (anonymous):

also which of the following quadratic functions has x=-1 and x=-2 as its only two real zeros A x^4-3x^3+3x^2-3x+2 B x^4+3x^3+3x^2-3x-2 C x ^4+3x^3+3x^2+3x+2 D x^4-3x^3-3x^2-3x-2

OpenStudy (anonymous):

See whether or not A, B, C, or D results in f(x)= 0. In other words, substitute each number in A,B,C and D for x until you find that both numbers in zero for f(x). If we test A, we see that \[f(1) = 1 - 6 +11 -6 =0 \] \[f(3) = 27 - 54 +33 -6 =0 \] Remember that we're given that f(3) = 0; so, try and find what the remain root is since you know f(1) = 0.

OpenStudy (anonymous):

**both numbers result in zero for f(x).

OpenStudy (anonymous):

Your other question is similar. Simply plug in each value you were given for x and see if it results in f(x) = 0.